# Spectrality of Prime Size Tiles

### v4

@@ 1 Introduction
Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$. It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [18, Corollary 2.4].
@@
The proof of the second theorem then reduces to proving any $p$ points in general linear positions are tiling, for which we first apply a group isomorphism to map them to the origin and $(1,0\ldots,0),(0,2,0,\ldots,0),\ldots,(0,\ldots,0,p-1)$, then we construct a tiling complement for such a set explicitly.
@@ 2 Preliminaries
is its orthogonal group in $\mathbb{Z}_{n}^{d}$. That(3) holds for cyclic subgroups is quite trivial, while non-cyclic subgroups can be viewed as products of cyclic ones, its orthogonal set is then the intersection over orthogonal sets of its cyclic factors.
@@ 3 Main results
Given any $n$ linearly indenpendent vectors $v_{1},\ldots,v_{n}$ on $\mathbb{Z}^{n}$, there exists a group automorphism of the additive group $\mathbb{Z}^{n}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,1,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,1)$. In particular, if $p$ is a prime number and n=p1n=p-1, then there exists a group automorphism of the additive group $\mathbb{Z}^{p-1}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,2,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,p-1)$.
Proof.
Let $V$ be the matrix whose columns are $v_{1},\ldots,v_{n}$, and $D$ the determinant of $V$. Clearly $D\neq 0$ because of linear independence.
Let $V^{\prime}$ be the adjugate matrix of $V$, i.e., if $U_{i,j}$ is the determinant of the matrix obtained by deleting the $i$-th row and the $j$-th column of $V$, then $V^{\prime}$ is the transpose of the matrix whose $i,j$-th entry is $(-1)^{i+j}U_{i,j}$. Then $V^{\prime}V=DI$, and thus $V^{\prime}$ is a group isomorphism from $\mathbb{Z}^{n}$ to $D\mathbb{Z}^{n}$ that maps $v_{1},\ldots,v_{p-1}$ to $(D,0,\ldots,0)$, $(0,D,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,D)$. The desired automorphism can then be constructed by composing the following two isomorphisms:
$\mathbb{Z}^{n}\xrightarrow{\qquad\qquad V^{\prime}\qquad\qquad}D\mathbb{Z}^{n}\xrightarrow{\qquad x\mapsto x/D\qquad}\mathbb{Z}^{n},$
If n=p1n=p-1, then $\operatorname{diag}(1,2,\ldots,{p-1})$ is also a group automorphism on $\mathbb{Z}_{p}^{p-1}$, thus the desired automorphism is constructed by a further composition:
$\mathbb{Z}^{p-1}\xrightarrow{\qquad V^{\prime}\qquad}D\mathbb{Z}^{p-1}\xrightarrow{\;x\mapsto x/D\;}\mathbb{Z}^{p-1}\xrightarrow{\;\operatorname{diag}(1,2,\ldots,{p-1})\otimes I\;}\mathbb{Z}^{p-1}\left(=\mathbb{Z}_{p}^{p-1}\oplus(p\mathbb{Z})^{p-1}\right),$
where $I$ is identity so that $\operatorname{diag}(1,2,\ldots,{p-1})$ only acts on the $\mathbb{Z}_{p}^{p-1}$ part. ∎
Lemma 7.
The set
$A=\{(0,0,\ldots,0),(1,0,\ldots,0),(0,2,0,\ldots,0),\ldots,(0,\ldots,0,p-1)\},$
@@
Let $x=(1,\ldots,1)$, and $\omega=e^{2\pi i/p}$, then it is easy to verify that
@@
Hence the cyclic group generated by $x$ (denoted by $\langle x\rangle$) is a spectrum of $A$. To see that $A$ is tiling, we will construct a set $B$ so that $|A|\cdot|B|=|\mathbb{Z}_{p}^{p-1}|$ and
@@
Since $p$ is a prime number, the (p1)(p-1)-st roots of unity in the finite field $\mathbb{F}_{p}$ is precisely $1,\ldots,p-1$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and the set of its columns, denoted by $u_{1},\ldots,u_{p-1}$, is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$). In particular, since $x$ is a column of it we may without loss of generality take $u_{1}=x$.
@@
Without loss of generality we may assume that one of the point is the origin, and the other p1p-1 points are $v_{1},\ldots,v_{p-1}$. By Lemma 6, we can find a group automorphism of $\mathbb{Z}^{d}$ that maps the original $p$ points $0,v_{1},\ldots,v_{p-1}$ to
$A=\{0,e_{1},2e_{2},\ldots,(p-1)e_{p-1}\},$
where $e_{i}$ is the $i$-th standard Euclidean basis. Since it is a group automorphism we assert that $A$ is tiling if and only the original $p$ points are tiling.
Now since $A$ is completely inside the subgroup spanned by $e_{1},\ldots,e_{p-1}$, it suffices for us to look at the case at d=p1d=p-1 only. Then by Lemma 7 we easily see that $A$ is indeed a tiling set in $\mathbb{Z}_{p}^{p-1}$, thus also in $\mathbb{Z}^{p-1}$. ∎
The condition that $p$ points have to be in general linear positions is tight in the sense that $3$ colinear points need not be tiling, for example, $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1, see also [19] for a characterization of prime power size tiles in $\mathbb{Z}$). The author does not know non-tiling examples for other values of $p$, but would expect that they exist for every pair $p,d$ with d<p1d<p-1.

### v5

@@ 1 Introduction
Here $p$ points being in general linear positions means that they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$. It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [18, Corollary 2.4].
@@
The proof of the second theorem then reduces to proving any $p$ points in general linear positions are tiling, for which we can simply show that they tile the sublattice they generate. This follows from the fact that the set formed by the origin and $(1,0\ldots,0)$, $(0,2,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,p-1)$ is tiling (a tiling complement can be found explicitly by computation), then we can map this set to the $p$ points we are looking at by a linear transform.
@@ 2 Preliminaries
is its orthogonal group in $\mathbb{Z}_{n}^{d}$. That (3) holds for cyclic subgroups is quite trivial, while non-cyclic subgroups can be viewed as products of cyclic ones, then its orthogonal set is simply the intersection over orthogonal sets of its cyclic factors.
@@ 3 Main results
Let $e_{i}$ be the $i$-th standard Euclidean basis, then the set
$A=\{0,e_{1},2e_{2},\ldots,(p-1)e_{p-1}\},$
@@
Let $x=(1,\ldots,1)$, and set $\omega=e^{2\pi i/p}$, then it is easy to verify that
@@
Therefore the cyclic group generated by $x$ (denoted by $\langle x\rangle$) is a spectrum of $A$. To see that $A$ is tiling, we will construct a set $B$ so that $|A|\cdot|B|=|\mathbb{Z}_{p}^{p-1}|$ and
@@
Since $p$ is a prime number, each of $1,\ldots,p-1$ is $a$ (p1)(p-1)-st root of unity in the finite field $\mathbb{F}_{p}$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and the set of its columns, denoted by $u_{1},\ldots,u_{p-1}$, is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$). In particular, since $x$ is a column of it we may without loss of generality order the columns in the way so that $u_{1}=x$.
@@
Corollary 1.
The set
$A=\{0,e_{1},\ldots,e_{p-1}\},$
is both tiling and spectral in $\mathbb{Z}^{p-1}$.
Proof.
It suffices to show that $A$ is tiling in $\mathbb{Z}_{p}^{p-1}$ since if $(A,B)$ is a tiling pair in $\mathbb{Z}_{p}^{p-1}$, then $(A,B\oplus(p\mathbb{Z})^{p-1})$ becomes a tiling pair in $\mathbb{Z}^{p-1}$.
Let
$S=\{0,e_{1},2e_{2},\ldots,(p-1)e_{p-1}\},$
be the set defined in Lemma 6, and set $D=\operatorname{diag}(1,2,\ldots,{p-1})$. Then $D$ is a group automorphism on $\mathbb{Z}_{p}^{p-1}$, its inverse is simply $D^{-1}=\operatorname{diag}(1,2^{-1},\ldots,(p-1)^{-1})$ where for each $k\in\{1,\ldots,p-1\}$, $k^{-1}$ is the multiplicative inverse of $k$ in $\mathbb{Z}_{p}$, each $k^{-1}$ exists since $p$ is a prime number.
By Lemma 6, $S$ is tiling in $\mathbb{Z}_{p}^{p-1}$, therefore $A=D^{-1}S$ is also tiling in $D^{-1}\mathbb{Z}_{p}^{p-1}$, but $D^{-1}\mathbb{Z}_{p}^{p-1}$ equals $\mathbb{Z}_{p}^{p-1}$, thus $A$ is indeed tiling in $\mathbb{Z}_{p}^{p-1}$. Its spectrality then follows from Theorem 1. ∎
@@
Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$. Let $V$ be the linear transform defined as
$Ve_{i}=\begin{cases}v_{i},&\text{if }i=1,\ldots,p-1;\\ e_{i},&\text{if }i>p-1.\end{cases}$
$V$ has full rank since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions. Further set
$A=\{0,v_{1},\ldots,v_{p-1}\},\qquad E=\{0,e_{1},\ldots,e_{p-1}\}.$
By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the space obtained by applying $V$ to $\operatorname{span}(E)$, which is a sublattice since $\operatorname{span}(E)=\mathbb{Z}^{p-1}$. This implies that $A$ is tiling in $\mathbb{Z}^{d}$, and its tiling complement can be produced explicitly: If $E^{\prime}$ is a tiling complement of $E$ in $\mathbb{Z}^{p-1}$, and $F$ is the set formed by taking one representative from each coset of $V\mathbb{Z}^{p-1}$ (viewed as an additive subgroup) in $\mathbb{Z}^{d}$, then $B=(VE^{\prime})\oplus F$ is a tiling complement of $A$. ∎
Theorem 2 produces some interesting consequences, for example we can assert that any $3$ points must be tiling as long as they are not on a line. On the other hand, this also shows that the condition that $p$ points have to be in general linear positions is tight, since $3$ colinear points indeed need not be tiling, e.g., $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1, see also [19] for a characterization of prime power size tiles in $\mathbb{Z}$).